# Boundary Constraint-free Biomechanical Model-Based Surface Matching for Intraoperative Liver Deformation Correction

### v1

@@ III-D Modifications for Unknown Zero Boundary Conditions and Force Locations
As mentioned in Section II-B, without boundary conditions, the stiffness matrix $\mathbf{K}$ is ill-conditioned and cannot be inverted to give a unique, stable solution. However, determining ZBCs and applied force locations is challenging during liver surgery.
@@
$\mathbf{K}\rightarrow\mathbf{K}+k_{ss}\mathbf{I},$ (16)
where $k_{ss}$ is the stiffness of the soft spring and $\mathbf{I}$ is the identity matrix. This is analogous to the damping term added to the stiffness matrix in the work presented by Peterlik et al. [3]. The added small stiffness increases the inertia of the liver model being deformed, preventing rigid motion without internal deformations. This approach enables us to allow the forces to be applied at any spatial location on the liver surface.
As a consequence of these two modifications, we no longer need to quantify the actual physical forces applied to the liver’s surface; rather, we empower the optimization algorithm to identify both the spatial distribution and magnitude of applied forces that optimally deform the preoperative model to match the intraoperative point cloud.
This novel formulation is a key contributor to our goal of developing an algorithm that allows us to register the preoperative and intraoperative without prior knowledge. As we shall show, these modifications can produce accurate registration results.
@@ IV-1 In silico phantom
As illustrated in Fig. 2, we simulated the deformation of the in silico phantom using a linear elastic model with Young’s modulus $E=1$ and a Poisson ratio $\nu=0.49$. Forces are applied perpendicular to the x-y plane on a liver mesh model obtained from OpenHELP [34]. The resulting deformed surface is cropped to generate the simulated intraoperative point cloud, consisting of 934 vertices. The volumetric preoperative mesh comprises 4291 vertices and 1919 tetrahedra.
@@ IV-2 In vitro phantoms
We created two silicone liver phantoms, A and B, each embedded with 53 and 176 fiducial markers, respectively, fabricated using a 3D-printed mold based on the liver model in OpenHELP [34]. To induce deformation, wedges featuring different gradients were strategically inserted underneath parts of the posterior side of the undeformed phantom, as depicted in Fig. 3 (b). The resulting phantom, with the wedge underneath, is referred to as the deformed phantom. We used the undeformed data as the preoperative data and the deformed data as the intraoperative data. The undeformed liver meshes, featuring around 9000 faces and 4000 vertices, were processed using TetGen [35] to generate the source volumetric meshes consisting of approximately 9000 volumetric vertices and 80000 tetrahedra.
Phantom A and Phantom B underwent three and two deformations respectively, generating five deformed phantoms. We number the deformed phantoms (No.1-3 from phantom A and No.4-5 from phantom B). CT scans were acquired for undeformed and deformed phantoms, and their surfaces and fiducial landmarks were manually segmented. Undeformed and deformed phantoms were manually aligned and refined with an ICP registration implemented by Open3D python library [36]. Initial registration errors are detailed in Table I.
Additionally, partial anterior surfaces were cropped to achieve surface visibility ratios ranging from approximately $18\%$ to $27\%$, calculated by the ratios of cropped surface areas to the full surface areas, as summarized in Table I.
@@ IV-3 Open-CAS dataset
: it contains three in silico phantoms and one in vitro phantom. We refer to [5] for more detail. Deformations of the in silico phantoms are created using a non-linear biomechanical model. Regarding the in vitro phantom, two intraoperative surfaces from a CT and stereo-endoscopic imaging are included.
@@ IV-4 Sparse dataset
We use Set044, Set057, Set067, and Set084 from the Sparse-data non-rigid registration challenge [22, 37], where partial ground truth is available. Intraoperative surfaces are presented with sparse 3D point clouds collected using an optically tracked stylus. We align the sparse data with the undeformed model as we did in the previous Phantoms dataset.
Evaluation. The registration error involves calculating nodal displacement errors for the simulated cases and Target Registration Error (TRE) for the phantom cases. Their calculations share a common form:
@@
Figure 4: Results of in silico phantom study. (a) Force recovery with correct boundary conditions (ZBCs) and force locations. (b) Examination of priors on registration results. (c) Optimization speed and accuracy through optimization step $\alpha$ update and Nesterov’s acceleration. (d) Effective prevention of overfitting to Gaussian noise ($\sigma$ = 3 mm) with L2 regularization.
@@
[…]
Figure 5: Qualitative comparison results of registration methods for Phantom datasets No.3 and No.5, presenting both initial undeformed and estimated deformed meshes yielded by all the methods. For cases with partial data, the edges of the complete deformed meshes are also displayed. The intraoperative point cloud is highlighted in red, and obvious disparities between the mesh and intraoperative point cloud are marked with arrows.
Implementation. The proposed method was implemented in C++ using the Eigen library [38], running on a 3.60 GHZ Intel i9-9900K CPU and 64 GB of RAM.
We use a linear elastic biomechanical liver model for all experiments. Since a linear biomechanical model is employed, Young’s modulus $E$ solely influences the scale of the estimated forces. For this study, we scale the force by setting $E$ to 1. Poisson coefficient is set as $\nu=0.49$, a typical value for modeling incompressible tissue [39]. Unless otherwise noted, the soft spring constant was set to $k_{ss}=0.01$, regularization weight to $\beta=0.05$, and the optimization algorithm was stopped at 200 iterations.
Methods for comparison. We use the official implementations of V2Snet[20] and GMM-FEM [17] as benchmarks for comparison. For V2Snet, we use the official weight included in the implementation. GMM-FEM [17] is an optimization-based method that incorporates the FEM as a strain energy term to regularize the CPD [18]. We implemented it using its suggested parameter values, which were tested in our in silico phantom described in Section V-A and gave the lowest errors. Furthermore, we include the Rigid Procrustes that use RANSAC ICP implemented in the Open3D library [36] with ground truth correspondences to estimate the lowest errors that may be achieved via the rigid registration.
@@ V-A In silico Phantom Validation
Next, we examine a more realistic situation where prior knowledge of the location of zero boundary conditions and forces is unknown. We eliminate ZBCs and set the soft spring constant $k_{ss}$ to 0.05. We explore two scenarios: the former assumes that the forces exist solely on the posterior surface, aligning with the assumption in [4]; the latter assumes that forces can be distributed across the entire surface. As illustrated in Fig. 4 (b), allowing the forces to be applied across the full liver surface yields superior registration results compared to confining the applied forces to the posterior surface. The results also indicate accurate registration results can be obtained without prior knowledge of force locations and ZBCs. As noted before, we are no longer computing the actual physical forces applied to the liver’s surface, but we rather empower the optimization algorithm to determine the fictitious forces that minimize the registration error. Subsequent experiments are conducted, assuming that forces can exist on the full surface.
@@
In the end, we examine the effect of the regularization term by introducing Gaussian noise with magnitudes $\sigma=3$ mm. As shown in Fig. 4 (d), the regularization term prevents the source model from fitting the noise as the iterations progress.
The computation depends on the number of FEM nodes and intraoperative point cloud, as our method relies on the FEM model and the close point operator. The average time per iteration for this dataset is 0.11 seconds, which leads to an overall computation time of 22 seconds over the 200 iterations.
@@ V-B In vitro Phantom Validation
Table I summarizes the results for the phantom dataset, comparing the whole and partial surface registration errors of our proposed method relative to those achieved by V2S and GMM-FEM. Our method features a mean TREs below 5 mm across all testing samples. Notably, the differences in mean TREs between our predictions for partial and full surfaces are negligible, typically less than 1 mm. In contrast, GMM-FEM tends to exhibit lower errors on full surfaces, as illustrated by the mean and max TRE on the full surface of No.5, which are 1.15 $\pm$ 1.62 (8.14) mm, compared to the partial surface, featuring an error of 17.54 $\pm$ 3.68 (24.10) mm. V2Snet encounters challenges in accurately predicting results for No.3, resulting in a mean TRE exceeding 14 mm and a max TRE larger than 30 mm.
Fig. 5 illustrates several qualitative results that provide the reader with a deeper and more tangible interpretation of the registration performance achieved by our proposed and comparative methods. Notably, blade-like structures are evident in the deformed meshes generated by V2Snet, as indicated by arrows in both No. 3 and No. 5. Specifically, for both the full and partial surface cases of the No.3 phantom, the estimated deformed mesh of V2Snet exhibits discrepancies in fully aligning with the ridge of the target surface. This misalignment is also noticeable in the estimation by GMM-FEM, particularly in the case of the partial surface for the No.3 phantom. In contrast, our method consistently produces reasonable deformations for the full and partial cases of the No.3 phantom. It’s crucial to note that in the partial surface case of No.3, the predicted left inferior ridge from all methods does not precisely match the inferior ridge of the full target surface. This discrepancy results from the inherent ambiguity associated with partially deformed surfaces, a phenomenon also observed in the partial surface of No.5.
@@ V-D Sparse dataset
TABLE III: Comparison of registration errors on the Sparse dataset. Mean ± Standard Deviation (Max) Error in millimeters is presented. The lowest mean values and the lowest maximum values are highlighted in bold.
Set Initial Rigid Procrustes GMM-FEM Proposed Method
044 11.15 $\pm$ 4.73 (24.10) 6.42 $\pm$ 5.11 (18.89) 5.69 $\pm$ 3.48 (16.68) 3.67 $\pm$ 1.24 (6.32)
057 14.25 $\pm$ 2.01 (19.39) 6.50 $\pm$ 5.34 (19.69) 4.43 $\pm$ 2.28 (10.33) 3.14 $\pm$ 1.34 (5.83)
067 8.08 $\pm$ 2.13 (13.58) 3.10 $\pm$ 2.35 (9.75) 6.35 $\pm$ 2.08 (10.28) 2.66 $\pm$ 1.41 (5.48)
084 14.74 $\pm$ 2.01 (19.39) 3.31 $\pm$ 2.57 (10.97) 4.43 $\pm$ 1.47 (7.29) 2.43 $\pm$ 1.11 (4.45)
The registration errors are detailed in Table III. Notably, V2Snet was excluded from this particular test as it necessitates intraoperative data in the form of meshes. Our method demonstrates consistent performance, with a mean TRE below 5 mm and a max TRE across all testing samples of less than 7 mm. In contrast, GMM-FEM exhibits a mean TRE exceeding 5 mm on Set044 and Set067, with a max TRE surpassing 7 mm in all samples.
@@
Qualitative results for Set044 in Fig. 7 reveal challenges for GMM-FEM in aligning the left and right inferior ridges of the liver, a task where our method excels in accurately fitting the intraoperative point cloud.
@@
Figure 7: Qualitative results for Set044 in the Sparse dataset, presenting the initial and estimated deformed meshes. The intraoperative point cloud is highlighted in red, and obvious disparities between the mesh and intraoperative point cloud are marked with arrows.

### v2

@@ III-D Modifications for Unknown Zero Boundary Conditions and Force Locations
As mentioned in Section II-B, without boundary conditions, the stiffness matrix $\mathbf{K}$ is ill-conditioned and cannot be inverted to yield a unique, stable solution. However, determining ZBCs and applied force locations is challenging during liver surgery.
@@
$\mathbf{K}\leftarrow\mathbf{K}+k_{ss}\mathbf{I},$ (12)
where $k_{ss}$ is the stiffness of the soft spring and $\mathbf{I}$ is the identity matrix. This is analogous to the work presented by Peterlik et al. [3], where a damping term was added to the stiffness matrix, which was then integrated into a strain/deformation energy term to regularize a data term. In our work, the damping term is integrated directly into the date term via the FEM. The added small stiffness increases the inertia of the liver model being deformed, preventing rigid motion without internal deformations. This approach enables us to allow the forces to be applied at any spatial location on the liver surface.
As a consequence of these modifications, we no longer need to quantify the actual location of ZBC’s and physical forces applied to the liver’s surface; rather, we empower the optimization algorithm to identify both the spatial distribution and magnitude of applied forces that optimally deform the preoperative model to match the intraoperative point cloud. This modified tangent stiffness matrix is incorporated into the gradient of the our data term described in Eq. 8.
This formulation is a key contributor to our goal of developing an algorithm that allows us to register the preoperative and intraoperative without prior knowledge. As we shall show, these modifications yield accurate registration results.
@@ IV-1 In silico phantom
As illustrated in Fig. 2, we simulated the deformation of the in silico phantom using a linear elastic model with Young’s modulus $E=1$ and a Poisson’s ratio $\nu=0.49$. Forces are applied perpendicular to the x-y plane on a liver mesh model obtained from OpenHELP [33]. The resulting deformed surface is cropped to generate the simulated intraoperative point cloud, consisting of 934 vertices. The volumetric preoperative mesh comprises 4291 vertices and 1919 tetrahedra.
@@
Figure 4: Distributions of fiducial markers in in vitro silicone phantoms A (left) and B (right).
@@ IV-2 In vitro phantoms
We created two liver phantoms, A and B, each embedded with 53 and 176 fiducial markers, respectively, as illustrated in Fig. 4, and fabricated using synthetic gelatin (Humimic Medical Gelatin #0) and a 3D-printed mold based on a liver model [34]. To induce deformation, wedges featuring different gradients were strategically inserted underneath parts of the posterior side of the undeformed phantom, as depicted in Fig. 3 (b). The resulting phantom, with the wedge underneath, is referred to as the deformed phantom. We used the undeformed data as the preoperative data and the deformed data as the intraoperative data. The undeformed liver meshes, featuring around 9000 faces and 4000 vertices, were processed using TetGen [35] to generate the source volumetric meshes consisting of approximately 9000 volumetric vertices and 80000 tetrahedra.
Phantom A and Phantom B underwent two deformations, respectively, generating four deformed phantom configurations. We number the deformed phantoms (No. 1-2 from phantom A and No. 3-4 from phantom B). CT scans of both undeformed and deformed phantoms were acquired, and their surfaces and fiducial landmarks were manually segmented. Undeformed and deformed phantoms were manually aligned and refined with an ICP registration implemented using the Open3D Python library [36]. Initial registration errors are detailed in Table I.
Additionally, partial anterior surfaces were cropped to achieve surface visibility ratios ranging from approximately $20\%$ to $27\%$, calculated as the ratios of cropped surface areas to the full surface areas, as summarized in Table I.
@@ IV-3 Open-CAS dataset
: The Open-CAS dataset contains three in silico phantoms and one in vitro phantom, as described in detailed in [5]. Deformations of the in silico phantoms are generated using a non-linear biomechanical model. The in vitro phantom featuring 6 fiducial markers was fabricated from silicone, and two intraoperative surfaces from a CT and stereo-endoscopic imaging are included.
@@ IV-4 Sparse dataset
We use the Sparse Non-rigid Registration Challenge [22, 37, 38], where 112 samples with ground truth are released [39]. Intraoperative surfaces are presented with sparse 3D point clouds collected using an optically tracked stylus. The dataset also provides initial rigid transformations to align the preoperative and intraoperative datasets. Lastly, the dataset also includes sub-dataset splits featuring different visibility ratios and noise magnitudes, as described in detail in [38].
Evaluation. The registration accuracy was estimated by calculating nodal displacement errors for the simulated cases and Target Registration Error (TRE) for the phantom cases. Their calculations share a common form:
@@
Figure 5: Results of in silico phantom study. Force recovery with correct boundary conditions (ZBCs) and force locations (left). Examination of priors on registration results (middle). Optimization speed and accuracy through optimization step $\alpha$ update and Nesterov’s acceleration (right).
@@
Figure 6: Qualitative comparison of the registration results yielded by different methods for Phantom datasets No. 2 and No. 4, presenting both initial undeformed and estimated deformed meshes yielded by all the methods. For cases with partial data, the edges of the complete deformed meshes are also displayed. The intraoperative point cloud is highlighted in red, and obvious disparities between the mesh and intraoperative point cloud are marked with arrows.
Implementation. The proposed method was implemented in C++ using the Eigen library [40], running on a 3.60 GHZ Intel i9-9900K CPU and 64 GB of RAM. The Cholesky decomposition implemented in the Eigen library, named SimplicialLLT, is used to invert the stiffness matrix from Eq. 12. An in-house implementation of a finite element model was used for the biomechanical calculations. We use a linear elastic biomechanical liver model for all experiments. Since a linear biomechanical model is employed, Young’s modulus $E$ solely influences the scale of the estimated forces and does not affect the displacement solution. For this study, we scale the force by setting $E$ to 1. Poisson’s coefficient is set as $\nu=0.49$, a typical value for modeling incompressible tissue [41]. The soft spring constant was set to $k_{ss}=0.01$, and the optimization algorithm was stopped at 200 iterations. The implemented procedure is detailed in Algorithm 1.
Open source methods for comparison. We use the official implementations of V2Snet[20] and GMM-FEM [17] as benchmarks for comparison. For V2Snet, we use the official weight included in the implementation. GMM-FEM [17] is an optimization-based method that incorporates the FEM as a strain energy term to regularize the CPD [18]. We implemented it using its suggested parameter values, which were tested in our in silico phantom described in Section V-A and gave the lowest errors. Furthermore, we include the Rigid Procrustes method that uses RANSAC ICP implemented in the Open3D library [36] with ground truth correspondences to estimate the lowest errors that may be achieved via the rigid registration.
Algorithm 1 Linear Elastic Surface Matching Registration
1:Inputs:
2:      x the preoperative (source) nodes of liver mesh
y the intraoperative (target) surface point cloud
$k_{ss}$ the soft spring constant
maxIters the maximum number of iterations
3:Outputs:
4:      u displacements for registered liver mesh
1:initialize displacements $\mathbf{u^{0}}\leftarrow 0$;
2:initialize forces $\mathbf{f^{0}}\leftarrow 0$;
3:calculate stiffness matrix $\mathbf{K}\leftarrow\mathbf{K}+k_{ss}$;
4:for k = 0 to maxIters do
5:     update mesh nodal displacements $\mathbf{u}^{k}=\mathbf{K}^{-1}\mathbf{f}^{k}$;
6:     update mesh nodes locations $\mathbf{x}\leftarrow\mathbf{x}+\mathbf{u}^{k}$;
7:     compute correspondence matrix C;
8:     update momentum term $\mathbf{p}^{k}$ (Eq.6);
9:     compute gradient of data term $\nabla J(\mathbf{p}^{k})$ (Eq.9);
10:     compute optimal step size $\alpha^{k}$ (Eq.11);
11:     update forces ${\mathbf{f}}^{k+1}$ (Eq.7);
12:end for
@@ V-A In Silico Phantom Validation
Next, we examine a more realistic situation where prior knowledge of the location of zero boundary conditions and forces is unknown. We eliminate ZBCs and include the soft spring constant $k_{ss}$. We explore two scenarios: the former assumes that the forces exist solely on the posterior surface, aligning with the assumption in [4]; the latter assumes that forces can be distributed across the entire surface. As illustrated in Fig. 5 (b), allowing the forces to be applied across the full liver surface yields superior registration results compared to confining the applied forces to the posterior surface. The results also indicate accurate registration results can be obtained without prior knowledge of force locations and ZBCs. As noted before, we are no longer computing the actual physical forces applied to the liver’s surface, but we rather empower the optimization algorithm to determine the fictitious forces that minimize the registration error. Subsequent experiments are conducted, assuming that forces can exist on the full surface.
@@
The computation depends on the number of FEM nodes and intraoperative point cloud, as our method relies on the FEM model and the closest point operator. The overall computation time for this dataset is 22 seconds over the 200 iterations.
@@ V-B In vitro Phantom Validation
Table I summarizes the results for the phantom dataset, comparing the whole and partial surface registration errors of our proposed method relative to those achieved by V2S and GMM-FEM. Our method features a mean TRE below 5 mm across all testing samples. Notably, the differences in mean TREs between our predictions for partial and full surfaces are negligible, typically less than 1 mm. In contrast, GMM-FEM tends to exhibit lower errors on full surfaces, as illustrated by the mean and max TRE on the full surface of No. 4, which are 1.15 $\pm$ 1.62 (8.14) mm, compared to the partial surface, featuring an error of 17.54 $\pm$ 3.68 (24.10) mm. V2Snet encounters challenges in accurately predicting results for No. 2, resulting in a mean TRE exceeding 5 mm and a max TRE larger than 15 mm.
Fig. 6 illustrates several qualitative results that provide the reader with a deeper and more tangible interpretation of the registration performance achieved by our proposed and comparative methods. Notably, blade-like structures are evident in the deformed meshes generated by V2Snet, as indicated by arrows in both No. 2 and No. 4. Specifically, for both the full and partial surface cases of the No. 2 phantom, the estimated deformed mesh of V2Snet exhibits discrepancies in fully aligning with the ridge of the target surface. This misalignment is also noticeable in the estimation by GMM-FEM, particularly in the case of the partial surface for the No. 2 phantom. In contrast, our method consistently produces reasonable deformations for both the full and partial cases of the No. 2 phantom.
@@ V-D Sparse Dataset
TABLE III: Comparison of registration errors on the Sparse dataset. Mean ± Standard Deviation (Median) of Mean Error in millimeters is presented. The lowest mean values and the lowest median values are highlighted in bold.
Method Type Visibility All
20-28% 28-36% 36-44%
V2S-net [20] Deep learning 7.79 $\pm$ 4.86 (6.61) 6.79 $\pm$ 4.63 (6.10) 8.24 $\pm$ 6.34 (6.86) 7.55 $\pm$ 5.28 (6.37)
Jia et al. [42] 4.22 $\pm$ 0.96 (4.02) 4.57 $\pm$ 2.79 (3.88) 4.02 $\pm$1.31 (3.72) 4.29 $\pm$ 1.93 (3.84)
Heiselman et al. [1] 3.27 $\pm$ 1.07 (2.89) 3.00 $\pm$ 0.67 (2.97) 2.99 $\pm$ 0.78 (2.73) 3.08 $\pm$ 0.85 (2.89)
Mestdagh et al. [16] 3.54 $\pm$ 1.11 (3.47) 3.27 $\pm$ 0.85 (3.19) 3.13 $\pm$ 0.82 (3.19) 3.31 $\pm$ 1.86 (2.95)
Ringel et al. [43] Biomechanical 4.56 $\pm$ 0.88 (4.48) 4.62 $\pm$ 0.99 (4.71) 4.62 $\pm$ 0.98 (4.37) 4.60 $\pm$ 0.95 (4.52)
AIC [44] model 5.06 $\pm$ 0.92 (4.93) 4.91 $\pm$ 1.11 (4.75) 4.61 $\pm$ 1.14 (4.52) 4.86 $\pm$ 1.07 (4.77)
GMM-FEM [17] (linear elasticity) 5.40 $\pm$ 1.56 (5.14) 5.14 $\pm$ 1.22 (4.83) 4.55 $\pm$ 1.16 (4.58) 5.00 $\pm$ 1.37 (4.75)
Proposed method 3.05 $\pm$ 0.75 (2.95) 2.94 $\pm$ 0.66 (2.74) 2.78 $\pm$ 0.68 (2.82) 2.93 $\pm$ 0.68 (2.82)
TABLE IV: Sensitivity analysis of the proposed method to the initial alignment analysis performed on the sparse dataset. Mean ± Standard Deviation (Median) of Mean Error in millimeters is presented.
Initial error w/ proposed method
Optimal rigid registration 3.77 $\pm$ 2.40 (3.16) 2.97 $\pm$ 0.66 (2.92)
Manually initialized ICP 5.53 $\pm$ 3.43 (4.84) 2.93 $\pm$ 0.68 (2.82)
wICP [27] 7.73 $\pm$ 4.48 (6.74) 2.97 $\pm$ 0.72 (2.85)
TABLE V: Performance of the proposed method to noise-free and noise-affected sparse datasets. Mean ± Standard Deviation of Mean Error in millimeters is presented.
Noise = 0 mm Noise = 2mm
Manually initialized ICP 5.53 $\pm$ 3.43 5.59 $\pm$ 1.61
w/ Heiselman et al. [1] 2.50 $\pm$ 0.31 3.34 $\pm$ 0.87
w/ Proposed method 2.23 $\pm$ 0.34 3.17 $\pm$ 0.66
Figure 8: Parameter sensitivity analysis for $k_{ss}$, $\nu$, and iteration number performed on the entire Sparse dataset. The median and mean values of mean errors are represented by orange lines and crosses, respectively.
The registration errors are detailed in Tables III, IV, and V. Except for GMM-FEM and our proposed method, results from other approaches align with previous reports [44, 38, 39]. The other methods are either closed-source [42, 1, 43], require additional data processing [20], or involve manual input [16]. The most similar method to our proposed method is that of Mestdagh et al. [16], which also uses a FEM as the data term but requires manual boundary condition identification, which is known to pose significant challenges. Lastly, all biomechanical-based methods employ the linear elastic model.
The sparse dataset includes three initial transformations for the rigid alignment. Table III and Table V report results using the manually initialized ICP transformation, consistent with the performance shown by Heiselsman et al. [1]. Table IV presents results from our proposed method using all initial transformations, demonstrating its robustness across different starting conditions. The proposed method consistently achieves the lowest mean registration errors across various settings, including the entire sparse dataset (Table III), different visibility conditions (Table III), multiple initial transformations (Table IV), and both noise-free and noise-affected scenarios (Table V), demonstrating its effectiveness and robustness. As shown in Table III, our proposed method outperforms the method developed by Mestdagh et al. [16], highlighting its effectiveness in automatically identifying boundary conditions. Compared to other FEM-based methods [1, 17], our approach also achieves lower registration errors (Table III and Table V). Lastly, the learning-based methods [42, 20] show less satisfactory performance than the best biomechanical approaches.
@@
Qualitative results for Set100 in Fig. 9 reveal challenges for GMM-FEM in aligning the right inferior ridges of the liver, where our method excels in accurately fitting the intraoperative point cloud.
@@
Figure 9: Qualitative results for Set0100 of the Sparse dataset, presenting the initial position from the manually initialized ICP, the official transformation, and estimated deformed meshes. The intraoperative point cloud is highlighted in red, and obvious disparities between the mesh and intraoperative point cloud are marked with arrows.